OpenStudy (anonymous):

how to rewrite the equation in function form. 7+7x=8

OpenStudy (anonymous):

are you sure you posted the question correctly?

OpenStudy (anonymous):

oops y+7x=8

OpenStudy (m):

y = -7x + 8

OpenStudy (m):

y = f(x) = -7x + 8

OpenStudy (anonymous):

8y-16x=4

OpenStudy (m):

y = (4 +16x)/8

OpenStudy (m):

you just want to isolate y

OpenStudy (anonymous):

ok, don;t understand the/8

OpenStudy (m):

it's a function becuase you can plug in for x and you'll get a function of x or a y

OpenStudy (m):

this is how you will plot points on xy plane graph

OpenStudy (m):

plug in for x coordinates and you get y coordinates

OpenStudy (m):

when they ask you to rearrange in function for then just solve for y lol

OpenStudy (anonymous):

tell wether the ordered pair is a solution of the equation? 7=3x+4;(1,6)

OpenStudy (m):

yes it is

OpenStudy (m):

when you see (1, 6) that usually means (x, y)

OpenStudy (m):

plug in 1 for x then 7 = 3 (1) + 4

OpenStudy (anonymous):

ok don't laugh at me heres another one, 7=-5x=2;(-3,-13)

OpenStudy (m):

you wrote it wrong i think

OpenStudy (anonymous):

y=-5x=2

OpenStudy (m):

no

OpenStudy (m):

-5(-3) = 15

OpenStudy (m):

y = -13 so it's wrong

OpenStudy (anonymous):

ok get it. how bout this 1. Your science class wants to visit an aquarium. Tickets cost $12.95/person. You also need to pay for the bus, which is $120. use the equation C=12.95p+120. where C is the total cost of the trip and p is the number of people going. If the final cost is $767.50, how many people went on the trip?

OpenStudy (m):

just plug in 767.50 for C

OpenStudy (m):

then solve for p

OpenStudy (m):

(767.50 - 120)/12.95 = p

OpenStudy (anonymous):

o ok, thank you sooooo much

OpenStudy (m):

p = 50

OpenStudy (m):

ok good luck :)